# Decomposing the Automorphism group of the surreal numbers

### v1

@@ 5. Preserving exponentiation and the $\omega$-map
Now we turn to automorphisms of $(\mathbf{No},+,\cdot,\exp)$. As an ordered exponential field $\mathbf{No}$ is strongly homogeneous: for any set-sized elementary exponential subfield $K\subseteq\mathbf{No}$ and any elementary embedding $\sigma\colon K\to\mathbf{No}$, there is an automorphism of $(\mathbf{No},+,\cdot,\exp)$ extending $\sigma$. In addition, any automorphism of $\mathbf{No}$ arising from a composition as in Remark 4.1 is an automorphism of $(\mathbf{No},+,\cdot,\exp)$. While automorphisms of $(\mathbf{No},+,\cdot,\exp)$ are plentiful, we have not been able to find an example of a nontrivial $1$-automorphism of $(\mathbf{No},+,\cdot,\exp)$. This is partially explained by the following negative result:
@@
Let $\partial\colon\mathbf{No}\to\mathbf{No}$ be a contractive strongly $\mathbb{R}$-linear derivation and suppose that $\partial$ is not identically zero. In light of 4.2, it is enough to show that the automorphism $\sigma\coloneqq\sum_{i=0}^{\infty}\frac{1}{i!}\partial^{i}$ does not commute with the exponential map.
@@
Let $\varepsilon\coloneqq\sigma(y)-y$, so $\exp(\sigma y)=\exp(y+\varepsilon)=\exp(y)\exp(\varepsilon)$. Since $\partial$ is contractive, we have $v(\varepsilon)=v(\partial y)<0$, so $v\exp(\varepsilon)\neq 0$. Therefore,
@@
where the last equality uses that $\sigma$ is a $1$-automorphism. Thus, $\sigma$ doesn’t commute with the exponential. ∎
Automorphisms obtained from compositions as in Remark 4.1 are always strongly $\mathbb{R}$-linear and commute with the exponential. Thus we have:
@@
Are there any nontrivial $1$-automorphisms of $(\mathbf{No},+,\cdot,\exp)$?
@@
What are the nontrivial automorphisms of $(\mathbf{No},+,\cdot,\omega^{x})$?
@@
Acknowledgements
This project was initiated within the Ontario/Baden-Württemberg Faculty Mobility Program in August 2023 and resumed at the Banff International Research Station in February 2024. We gratefully acknowledge the involved institution for partial funding and their hospitality. We also thank Salma Kuhlmann for suggesting this project, Patrick Speissegger for his support, and Vincent Bagayoko for helpful discussions. The first author was supported in part by the National Science Foundation under Award No. DMS-2103240. Part of this research were conducted while the first author was hosted by the Max Planck Institute for Mathematics, and he thanks the MPIM for its support and hospitality. The second author was partially supported by the Vector Stiftung.

### v2

@@ 5. Preserving exponentiation and the $\omega$-map
Now we turn to automorphisms of $(\mathbf{No},+,\cdot,\exp)$. As an ordered exponential field $\mathbf{No}$ is strongly homogeneous: for any set-sized elementary exponential subfield $K\subseteq\mathbf{No}$ and any elementary embedding $\sigma\colon K\to\mathbf{No}$, there is an automorphism of $(\mathbf{No},+,\cdot,\exp)$ extending $\sigma$. While automorphisms of $(\mathbf{No},+,\cdot,\exp)$ are plentiful, we have not been able to find an example of a non-trivial $1$-automorphism of $(\mathbf{No},+,\cdot,\exp)$. This is partially explained by the following negative result.
@@
Let $\partial\colon\mathbf{No}\to\mathbf{No}$ be a contracting strongly $\mathbb{R}$-linear derivation and suppose that $\partial$ is not identically zero. In light of 4.2, it is enough to show that the automorphism $\sigma\coloneqq\sum_{i=0}^{\infty}\frac{1}{i!}\partial^{i}$ does not commute with the exponential map.
@@
Now with $y$ as above, let $\varepsilon\coloneqq\sigma(y)-y$, so $\exp(\sigma y)=\exp(y+\varepsilon)=\exp(y)\exp(\varepsilon)$. Since $\partial$ is contracting, we have $v(\varepsilon)=v(\partial y)<0$, so $v\exp(\varepsilon)\neq 0$. Therefore,
@@
where the last equality uses that $\sigma$ is a $1$-automorphism. Thus, $\sigma$ does not commute with the exponential. ∎
Automorphisms obtained from compositions as in Remark 4.1 are always strongly $\mathbb{R}$-linear and commute with the exponential. Thus we have the following.
@@
Are there any non-trivial $1$-automorphisms of $(\mathbf{No},+,\cdot,\exp)$?
@@
What are the non-trivial automorphisms of $(\mathbf{No},+,\cdot,\omega^{x})$?
@@
Acknowledgments
This project was initiated within the Ontario/Baden-Württemberg Faculty Mobility Program in August 2023 and resumed at the Banff International Research Station in February 2024. We gratefully acknowledge the involved institutions for partial funding and their hospitality. The first author is a postdoctoral researcher of the Fonds de la Recherche Scientifique – FNRS, and was supported in part by the National Science Foundation under Award No. DMS-2103240. Parts of this research were conducted while the first author was hosted by the Max Planck Institute for Mathematics, and he thanks the MPIM for its support and hospitality. The second author was partially supported by the Vector Stiftung.
We also thank Salma Kuhlmann for suggesting this project, Patrick Speissegger for his support, Vincent Bagayoko for helpful discussions and Mickaël Matusinski for pointing out Question 4.5.